OpenStudy (anonymous):

f(x)=√5-5x when x=-2,8

OpenStudy (anonymous):

Just plug in -2 for x to find the value, and then plug in 8 for x

OpenStudy (anonymous):

I tried that and the answer was wrong

OpenStudy (anonymous):

are u still there?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

You probably plugged it in wrong

OpenStudy (anonymous):

What did you get when you plugged in -2

OpenStudy (anonymous):

The answer I got was √-35

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

f(8)=√5-5(8)

OpenStudy (anonymous):

I do not know! this is why I am confused!

OpenStudy (anonymous):

Is it \[\sqrt{5}-5x \] or \[\sqrt{5-5x}\]

OpenStudy (anonymous):

The 2nd one. I guess I cant find that symbol!

OpenStudy (anonymous):

Ok, when x=-2, y=\[\sqrt{15}\]

OpenStudy (anonymous):

ok, can u show me how you got that please?